Grandmother is 4 times as old as grandson 5 years ago. How old is the grandmother if their present age is 55 years together?
Please do it in units not with x and y variables.

Answer by ikleyn(52786)

.

grandson + 4*(grandson-5) = 55


5*grandson = 55 + 20 = 75


grandson = 75/5 = 15 years old.    ANSWER